Express as logarithms of single number 2-2log 5 base 3
Answer by Theo(13342)

your expression is 2 minus 2 * log3(5).

2 is equivalent to log3(9).

your expression becomes log3(9) - log3(25)

this becomes log3(9/25)

log3(9/25) is equal to log(9/25)/log(3) = -.9299470414

2 - 2 * log3(5) is equivalent to 2 - 2 * log(5)/log(3) which is equal to -.9299470414.

they give you the same answer, so 2 - 2 * log3(5) is equivalent to log3(9/25).

a couple of basic properties of logs are active here.

these are the properties as they apply to this problem.

log3(x) = log(x) / log(3)

log3(x) = y if and only if 3^y = x


your solution is that 2 - 2 * log3(5) is equivalent to log3(9/25).